Shift report 20010731 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jul 31 2001 - 23:59:48 EDT

  • Next message: Apache: "Shift report 20010801 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jul 31 23:59:47 2001, Shift: 16:00-24:00
    Supervisor: KH, with: 
    Log book#6 pages 133 - 137
    
    Shift Summary:
    ------------------------------------------------------------
    Had about 3 hr 10 min of beam during this shift
    
    Run 3948 16:00-16:35  28325 collisions (in progress at start)
    Run 3951 19:25-20:35   4525 collisions
    Run 3952 22:30-24:00 ~30000 collisions at shift end; continuing
    
    The run in progress started with a count rate of ~20 events/sec.
    
    
    Machine status:
    Starting at 16:35 they did some studies on RF, then dumped 
    the beam and reinjected for the run that started at 19:25
    About 20:35, there was a quench event which about 2 hrs
    to recover from.
    
    Summary of collisions from run 3928 on
    run     #collisions        comment
    3928     41943
    3932     21366
    3934     35242
    3935     24267
    3937      3316
    3941     89008        D1,D2 turned on in this run
    3944      6312        TOFW found off during this run
    3947      1845
    3948     28325        RICH turned on starting this run
    3951      4527
    3952     27717 estimated at shift end
    
    total   283569
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 01 2001 - 00:00:51 EDT